# Geometry and Optimization of Shallow Polynomial Networks

### v1

@@ 5 Conclusions
We investigated various aspects of shallow networks with polynomial activations. We described the functional space of these models and explored the relationship between network width and the loss landscape. Additionally, we examined teacher-student problems, relating them to low-rank tensor approximations and showing how certain discriminants encode the relation between data and optimization. We also focused on networks with quadratic activations, presenting variations of the Eckart-Young Theorem for teacher-student problems under different norms.
@@ Acknowledgments.
J.K. and E.P. were supported partially by NSF DMS 2309782, NSF CISE-IIS 2312746 and DE SC0025312, as well as start-up grants and a JTO fellowship from the University of Texas at Austin. YA was supported by the Israel Science Foundation (grant No. 724/22).
Yossi Arjevani: The Hebrew University. (yossi.arjevani@gmail.com)
Joan Bruna: New York University. (bruna@cims.nyu.edu)
Joe Kileel: University of Texas at Austin. (jkilee@math.utexas.edu)
Elzbieta Polak: University of Texas at Austin. (epolak@utexas.edu)
@@
Matthew Trager: AWS AI Labs (work done outside of Amazon). (matthewtrager@gmail.com)
Appendix A Proofs and Additional Results
@@
$\tau_{r}^{-1}(S)=\{(\alpha,W)\colon W^{T}diag(\alpha)W=S\}\subset\mathbb{R}^{r}\times\mathbb{R}^{r\times n},$
@@
    if ${\rm rk}(S)=r$ and $s_{+}>0$ $s_{-}>0$, then $\tau_{r}^{-1}(S)$ has $4\cdot\binom{r}{s_{+}}=4\cdot\binom{r}{s_{-}}$ connected components;
@@
It is enough to observe that $W^{T}diag(\alpha)W$ can have at most $\alpha_{+}$ strictly positive eigenvalues. Indeed, $U=\{v\in\mathbb{R}^{n}\,\,\colon\,\,(Wv)_{i}=0,\,\,\forall i\,\,s.t.\,\,\alpha_{i}>0\}$ is a vector space of dimension at least $n-\alpha_{+}$ where the restriction of the quadratic form $v\mapsto v^{T}W^{T}diag(\alpha)Wv$ is negative semi-definite. The same argument shows that $W^{T}diag(\alpha)W$ can have at most $\alpha_{-}$ strictly negative eigenvalues. ∎
@@
If ${\rm rk}(S)>r$ then $\tau_{r}^{-1}(S)$ is empty since any matrix of the form $W^{T}diag(\alpha)W$ has rank at most $r$. Assume now ${\rm rk}(S)=r$. If $(\alpha,W)\in\tau_{r}^{-1}(S)$, then all elements of $\alpha\in\mathbb{R}^{r}$ are non-zero, and by Lemma 38 exacly $s_{+}$ elements are positive and $s_{-}$ elements are negative. Equivalently, if we decompose the parameter space based on the sign of the elements of $\alpha$:
@@
then $\tau_{r}^{-1}(S)\cap\mathcal{U}_{\sigma}$ is not empty if and only if |$|\{i\colon\sigma_{i}=1\}|=s_{+}$, |$|\{i\colon\sigma_{i}=-1\}|=s_{-}$, and |$|\{i\colon\sigma_{i}=0\}|=0$. This yields $\binom{r}{s_{+}}=\binom{r}{s_{-}}$ vectors $\sigma$ such that $\tau_{r}^{-1}(S)\cap\mathcal{U}_{\sigma}$ is not empty. Because $\mathcal{U}_{\sigma}$ with $\sigma_{i}\neq 0$ are disconnected, the connected components of $\tau_{r}^{-1}(S)$ are the union of the connected components of each $\tau_{r}^{-1}(S)\cap\mathcal{U}_{\sigma}$. We can thus describe the connected components of the fiber $\tau_{r}^{-1}(S)$ with the constraint that the signs of $\alpha_{i}$ are fixed. It is easy to see that we can in fact further assume that $\alpha_{i}=\pm 1$, since there is always a continuous path within the fiber $\tau_{r}^{-1}(S)$ that “normalizes” each $\alpha_{i}$:
$\alpha_{i}(t)=(1-t)\alpha_{i}+t\frac{\alpha_{i}}{|\alpha_{i}|},\quad w_{i}(t)=\frac{w_{i}}{\sqrt{|\alpha_{i}(t)|}},\quad W(t)^{T}diag(\alpha(t))W(t)=S.$
@@
Writing $\mathcal{V}_{\sigma}=\{(\alpha,W)\colon\alpha_{i}=\sigma_{i}\}$ (so $\mathcal{V}_{\sigma}\subset\mathcal{U}_{\sigma}$), we focus on the connected components of $\tau_{r}^{-1}(S)\cap\mathcal{V}_{\sigma}$(formally, $\tau_{r}^{-1}(S)\cap\mathcal{V}_{\sigma}$ is a deformation retract of $\tau_{r}^{-1}(S)\cap\mathcal{U}_{\sigma}$, so it has the same number of connected components). By symmetry, we may assume that the first $s_{+}$ elments of $\alpha$ are $1$ and the last $s_{-}$ elements are $-1$. We consider the set
$O(s_{+},s_{-})=\{M\colon M^{T}diag(\sigma)M=diag(\sigma)\}\subset\mathbb{R}^{n\times n}.$ (34)
@@
If $s_{+}=0$ or $s_{-}=0$, this is the orthogonal group, otherwise it is the “indefinite” orthgonal group. The group $O(s_{+},s_{-})$ acts on $\tau_{r}^{-1}(S)\cap\mathcal{V}_{\sigma}$ according to $(M,(\sigma,W))\mapsto(\sigma,MW)$ since $W^{T}M^{T}diag(\sigma)MW=W^{T}diag(\sigma)W$. We claim that this action is transitive (any point in $\tau_{r}^{-1}(S)\cap\mathcal{V}_{\sigma}$ can be mapped to any other) and free (only the identity fixes any point). Indeed, if $W$ and $W^{\prime}$ are matrices in $\mathbb{R}^{r\times n}$ such that $W^{T}diag(\sigma)W=W^{\prime T}diag(\sigma)W^{\prime}=S$, then $Row(W)=Row(W^{\prime})=Row(S)$, and there exists $M\in GL(r,\mathbb{R})$ such that $W^{\prime}=MW$. Such $M$ belongs to $O(s_{+},s_{-})$ since $W^{T}M^{T}diag(\sigma)MW=W^{T}diag(\sigma)W$ implies $M^{T}diag(\sigma)M=Id_{r}$ (by multiplying by the pseudoiverses of $W^{T}$ and $W$). This proves that the action is transitive. The action is also free since if $(\sigma,W)\in\tau_{r}^{-1}(S)$ then $rk(W)=r$ and $MW=W$ implies $M=Id$. The fact that the action is transitive and free means that the number of connected components of $\tau_{r}^{-1}(S)\cap\mathcal{V}_{\sigma}$ is the same as the number of connected components $O(s_{+},s_{-})$. It is known that $O(s_{+},s_{-})$ has two connected components if $s_{+}=0$ or $s_{-}=0$, and four connected components otherwise [31]. In the former case, the two connected components correspond to the sign of the determinant, while in the latter case the components arise from the combinations of signs of the determinants of the two diagonal blocks of sizes $s_{+}\times s_{+}$ and $s_{-}\times s_{-}$. This proves cases $2$ and $3$.
@@
We now assume that $rk(S)<r$. We say that $S=W^{T}diag(\alpha)W$ is a “minimal factorization” of $S$ if $\alpha$ has exactly ${\rm rk}(S)=s_{+}+s_{-}$ non-zero elements that are all $1$ or $-1$ and if $w_{i}=0$ for all $i$ such that $\alpha_{i}=0$. In this setting, $\alpha$ must have exactly $s_{+}$ elements that are $1$ and $s_{-}$ elements that are $-1$ (Lemma 38). We will first prove that for every element $(\alpha,W)\in\tau_{r}^{-1}(S)$, we can find a path inside $\tau_{r}^{-1}(S)$ from $(\alpha,W)$ to a minimal factorization. Then we will show that any two minimal factorizations of $S$ are path connected $\tau_{r}^{-1}(S)$.
@@
For the first claim, we observe as before that it is always possible to rescale $\alpha_{i}$ so that all of its elements are $1,-1$ or 00. We thus need to show that we may transform $\alpha$ so it has exactly ${\rm rk}(S)$ non-zero elements. We prove this inductively arguing that if ${\rm rk}(S)<r$ and $W^{T}diag(\alpha)W=S$, we can find a path inside $\tau_{r}^{-1}(S)$ from $(\alpha,W)$ to $(\alpha^{\prime},W^{\prime})$ so that $\alpha_{i}^{\prime}=w_{i}^{\prime}=0$ for some $i$. The same argument can then be repeated for the remaining $r^{\prime}=r-1$ vectors $w_{i}$ and elements $\alpha_{i}$ and iterated until $r^{\prime}={\rm rk}(S)$. If $\alpha_{i}=0$, then we can simply scale the corresponding $w_{i}$ and, similarly, if $w_{i}=0$ we can scale the corresponding $\alpha_{i}$. If $\alpha_{i}\neq 0$ and $w_{i}\neq 0$ for all $i$, then necessarily ${\rm rk}(W)<r$, so there exists $v\in\mathbb{R}^{r}$ such that $v^{T}W=0$. Assume that $\alpha$ has the first $\alpha_{+}$ elements equal to $1$, and the last $\alpha_{-}$ elements equal to $-1$ (with $\alpha_{+}+\alpha_{-}=r$). If $\alpha_{+}=0$ or $\alpha_{-}=0$, then we consider a path $(\alpha(t),W(t))$ where $\alpha(t)=\alpha$, $W(t)=M(t)W$ and $M(t)$ is a path in $SO(r)$ such that $M(0)=Id_{r}$ and the first row of $M(1)$ is equal to $v/\|v\|$ where $v^{T}W=0$ (this path is exists because $SO(r)$ is connected). By construction the first row $M(1)W$ is zero, and we the can rescale the corresponding $\alpha_{1}$ to zero, as desired. If $\alpha_{+}>0$ and $\alpha_{-}>0$ then we apply essentially the same argument using the connected components of the indefinite orthogonal group $O(\alpha_{+},\alpha_{-})=\{M\in\mathbb{R}^{r\times r}\colon M^{T}diag(\alpha)M=diag(\alpha)\}$. However, we must first consider the special situation that our chosen $v\in\mathbb{R}^{k}\setminus\{0\}$ that satisfies $v^{T}W=0$ is such that $v^{T}diag(\alpha)v=0$ (i.e., it is an isotropic vector). If this is the case, we construct a path as follows. Letting $y=diag(\alpha)v$ and assuming that $y_{i}$ is a non-zero element of $y$ (note $y\neq 0$), we set $W(t)=W-(t/y_{i})yw_{i}^{T}$. Observing that $diag(\alpha)y=v$ and that $y^{T}diag(\alpha)y=v^{T}diag(\alpha)^{3}v=0$ (because $diag(\alpha)diag(\alpha)=Id$), we have that
$(W-(t/y_{i})yw_{i}^{T})^{T}diag(\alpha)(W-(t/y_{i})yw_{i}^{T})=W^{T}diag(\alpha)W,$
@@
and the $i$-th row of $W-(1/y_{i})yw_{i}^{T}$ is 00, as desired. We finally assume that $v\in\mathbb{R}^{k}$ for which $v^{T}W=0$ is such that $v^{T}diag(\alpha)v\neq 0$. We can normalize $v$ such that $v^{T}diag(\alpha)v=1$ or $v^{T}diag(\alpha)v=-1$. If $v^{T}diag(\alpha)v=1$, then we can find a path $M(t)$ in $O(\alpha_{+},\alpha_{-})$ such that $M(0)=Id_{r}$ and the first row of $M(1)$ is equal to $v$ or $-v$ (this is true in light of the description of the connected components of $O(\alpha_{+},\alpha_{-})$). [^4] It is helpful to note that $M\in O(\alpha_{+},\alpha_{-})$ if and only if $M^{T}\in O(\alpha_{+},\alpha_{-})$. Indeed: $M^{T}diag(\alpha)M=diag(\alpha)\Leftrightarrow M^{T}diag(\alpha)Mdiag(\alpha)M^{T}=M^{T}\Leftrightarrow Mdiag(\alpha)M^{T}=diag(\alpha)$. Similarly, if $v^{T}diag(\alpha)v=-1$, we can find a path $M(t)$ in $O(\alpha_{+},\alpha_{-})$ such that $M(0)=Id_{r}$ and the $(\alpha_{+})+1$-th row of $M(1)$ is equal to $v$ or $-v$. Either way, we have that $(\alpha,W(t))=(\alpha,M(t)W)$ is a path in $\tau_{r}^{-1}(S)$ such that $W(0)=W$ and and $W(1)$ has a zero row. This concludes the proof of the existence of a path in $\tau_{r}^{-1}(S)$ from any $(\alpha,W)$ to a minimal factorization of $S$.

### v2

@@ 5 Conclusions
We investigated various aspects of shallow networks with polynomial activations. We described the function space of these models and explored the relationship between network width and the loss landscape. Additionally, we examined teacher-student problems, relating them to low-rank tensor approximations and showing how certain discriminants encode the relation between data and optimization. We also focused on networks with quadratic activations, presenting variations of the Eckart-Young Theorem for teacher-student problems under different norms.
@@ Acknowledgments.
J.K. and E.P. were supported partially by NSF DMS 2309782, NSF CISE-IIS 2312746 and DE SC0025312, as well as start-up grants and a JTO fellowship from the University of Texas at Austin. YA was partially supported by the Israel Science Foundation (grant No. 724/22).
Yossi Arjevani: The Hebrew University (yossi.arjevani@gmail.com)
Joan Bruna: New York University (bruna@cims.nyu.edu)
Joe Kileel: University of Texas at Austin (jkileel@math.utexas.edu)
Elzbieta Polak: University of Texas at Austin (epolak@utexas.edu)
@@
Matthew Trager: AWS AI Labs, work done outside of Amazon (matthewtrager@gmail.com)
Appendix A Numerical Experiment
We present a numerical experiment to illustrate the results of Section 4. We fix a teacher quadratic model $\mathcal{V}=(\beta,V)\in\mathbb{R}^{5}\times\mathbb{R}^{5\times 5}$ defined by
$\beta=\begin{bmatrix}-4&-2&1&3&5\end{bmatrix},\qquad V=I-\tfrac{2}{5}J,$
where $I\in\mathbb{R}^{5\times 5}$ is the identity matrix and $J\in\mathbb{R}^{5\times 5}$ is the all-ones matrix. Since $V$ is orthogonal, the eigenvalues of the teacher matrix $T=V^{\top}\operatorname{diag}(\beta)V\in\operatorname{Sym}^{2}(\mathbb{R}^{5})$ are precisely $\beta$. We generate a dataset $\{(x_{i},y_{i})\}_{i=1}^{N}\subset\mathbb{R}^{5}\times\mathbb{R}$ of size $N=50,000$, where inputs $x_{i}\in\mathbb{R}^{5}$ are sampled from a standard Gaussian distribution, and outputs are computed as $y_{i}=f_{\mathcal{V}}(x_{i})=x_{i}^{\top}Tx_{i}$. This dataset is used to train a student model $f_{\mathcal{W}}$ of width $r=3$ by minimizing the mean squared error loss $L(\mathcal{W})=\frac{1}{N}\sum_{i=1}^{N}|y_{i}-f_{\mathcal{W}}(x_{i})|^{2}$. We use stochastic gradient descent (SGD) with a batch size of 256, a learning rate of $10^{-4}$, and train for 500 epochs, with Pytorch’s default initialization of linear layers. The experiment is repeated 1000 times, fixing the teacher model $\mathcal{V}$ across all trials.
From Corollary 32, the critical points in parameter space correspond to the union of the critical points in function space of the functional loss restricted to the smooth loci of the low-rank manifolds $\mathcal{S}(r,5)$ for $r\leq 3$, i.e. $\tau_{r}(\operatorname{Crit}(L))=\bigcup_{r\leq 3}\operatorname{Crit}(h_{T}|_{\mathcal{S}(r,5)})$. Here, the functional loss is given by the empirical risk squared norm $h_{T}(S)=\|T-S\|^{2}_{ERM}$ (12), which in the large-sample limit converges to the Gaussian norm $h_{T}(S)=2\langle S-T,S-T\rangle_{F}+\operatorname{tr}(S-T)^{2}$. Within each manifold $\mathcal{S}(r,5)$, all critical points and their indices are then completely described by Theorem 36. In our case, this results in a total of $\binom{5}{3}+\binom{5}{2}+\binom{5}{1}=25$ non-zero critical points. Computing the indices of these critical points using (31), we find that there are $2+2+2=6$ local minima across the manifolds $\mathcal{S}(3,5)$, $\mathcal{S}(2,5)$, and $\mathcal{S}(1,5)$, respectively. Note, however, that these low-rank minima are not necessarily minima in parameter space, though any minimum in parameter space must correspond to one of them. For example, in the notation of Theorem 36, the global minimum corresponds to the index set $\mathcal{I}=\{1,4,5\}$, which defines the student matrix $S_{\mathcal{I}}=V^{\top}{\rm diag}(\gamma_{\mathcal{I}})V$, where
$\gamma_{\mathcal{I}}=\begin{bmatrix}-4.2&0&0&2.8&4.8\end{bmatrix}$.
The values of $\gamma_{\mathcal{I}}$ are computed by adding $c_{\mathcal{I}}=\frac{-2+1}{5}=-0.2$ to the selected eigenvalues of the teacher matrix.
Figure 3(a) shows the eigenvalue histograms of the student models at the end of training, which align closely with the eigenvalues of the theoretically predicted minima. The average Frobenius distance between trained students and their nearest theoretical minima was approximately $8\times 10^{-2}$. We also performed direct optimization of the Gaussian norm (without sampling data) with a learning rate of $10^{-3}$ and 10,000 iterations. The resulting eigenvalue distributions (Figure 3(b)) again match the predicted minima and agree closely with the sample-based results. Figure 3(c) shows the frequency with which each theoretical minimum is closest to the trained student models. Only 4 of the 6 minima were actually reached in our trials. Finally, we repeated the experiment, replacing the standard Gaussian distribution used to generate inputs with a coordinate-wise uniform distribution in ([-1, 1]). Using the exact same training parameters, the final student models were significantly more dispersed, in line with Theorem 37: with a tolerance of $0.05$ Frobenius norm, there were 42 distinct students at the end of training, compared to only 5 in the Gaussian setting.
(a) Eigenvalue histogram for final students trained with sample-based mean squared error loss.
(b) Eigenvalue histogram for final students trained with Gaussian square norm minimization.
(c) Counts of nearest minima reached by student models, where ‘Min 6’ corresponds to the global minimum.
Figure 3: (a)–(b) Eigenvalue histograms of student models after training, overlaid with the predicted minima. Both sample-based and norm-based optimization yield nearly identical spectra that closely match the eigenvalues of the predicted minima. (c) Frequencies of minima reached (sample-based setting; results are nearly identical for the norm-based setting)
Appendix B Proofs and Additional Results
@@
$\tau_{r}^{-1}(S)=\{(\alpha,W)\colon W^{T}\operatorname{diag}(\alpha)W=S\}\subset\mathbb{R}^{r}\times\mathbb{R}^{r\times n},$
@@
    if ${\rm rk}(S)=r$, $s_{+}>0$, and $s_{-}>0$, then $\tau_{r}^{-1}(S)$ has $4\cdot\binom{r}{s_{+}}=4\cdot\binom{r}{s_{-}}$ connected components;
@@
It is enough to observe that $W^{T}\operatorname{diag}(\alpha)W$ can have at most $\alpha_{+}$ strictly positive eigenvalues. Indeed, $U=\{v\in\mathbb{R}^{n}\,\,\colon\,\,(Wv)_{i}=0,\,\,\forall i\,\,s.t.\,\,\alpha_{i}>0\}$ is a vector space of dimension at least $n-\alpha_{+}$ where the restriction of the quadratic form $v\mapsto v^{T}W^{T}\operatorname{diag}(\alpha)Wv$ is negative semi-definite. The same argument shows that $W^{T}\operatorname{diag}(\alpha)W$ can have at most $\alpha_{-}$ strictly negative eigenvalues. ∎
@@
If ${\rm rk}(S)>r$ then $\tau_{r}^{-1}(S)$ is empty since any matrix of the form $W^{T}\operatorname{diag}(\alpha)W$ has rank at most $r$. Assume now ${\rm rk}(S)=r$. If $(\alpha,W)\in\tau_{r}^{-1}(S)$, then all elements of $\alpha\in\mathbb{R}^{r}$ are non-zero, and by Lemma 39 exacly $s_{+}$ elements are positive and $s_{-}$ elements are negative. Equivalently, if we decompose the parameter space based on the sign of the elements of $\alpha$:
@@
then $\tau_{r}^{-1}(S)\cap\mathcal{U}_{\sigma}$ is not empty if and only if $|\{i\colon\sigma_{i}=1\}|=s_{+}$, $|\{i\colon\sigma_{i}=-1\}|=s_{-}$, and $|\{i\colon\sigma_{i}=0\}|=0$. This yields $\binom{r}{s_{+}}=\binom{r}{s_{-}}$ vectors $\sigma$ such that $\tau_{r}^{-1}(S)\cap\mathcal{U}_{\sigma}$ is not empty. Because $\mathcal{U}_{\sigma}$ with $\sigma_{i}\neq 0$ are disconnected, the connected components of $\tau_{r}^{-1}(S)$ are the union of the connected components of each $\tau_{r}^{-1}(S)\cap\mathcal{U}_{\sigma}$. We can thus describe the connected components of the fiber $\tau_{r}^{-1}(S)$ with the constraint that the signs of $\alpha_{i}$ are fixed. It is easy to see that we can in fact further assume that $\alpha_{i}=\pm 1$, since there is always a continuous path within the fiber $\tau_{r}^{-1}(S)$ that “normalizes” each $\alpha_{i}$:
$\alpha_{i}(t)=(1-t)\alpha_{i}+t\frac{\alpha_{i}}{|\alpha_{i}|},\quad w_{i}(t)=\frac{w_{i}}{\sqrt{|\alpha_{i}(t)|}},\quad W(t)^{T}\operatorname{diag}(\alpha(t))W(t)=S.$
@@
Writing $\mathcal{V}_{\sigma}=\{(\alpha,W)\colon\alpha_{i}=\sigma_{i}\}$ (so $\mathcal{V}_{\sigma}\subset\mathcal{U}_{\sigma}$), we focus on the connected components of $\tau_{r}^{-1}(S)\cap\mathcal{V}_{\sigma}$(formally, $\tau_{r}^{-1}(S)\cap\mathcal{V}_{\sigma}$ is a deformation retract of $\tau_{r}^{-1}(S)\cap\mathcal{U}_{\sigma}$, so it has the same number of connected components). By symmetry, we may assume that the first $s_{+}$ elments of $\alpha$ are 11 and the last $s_{-}$ elements are 1-1. We consider the set
$O(s_{+},s_{-})=\{M\colon M^{T}\operatorname{diag}(\sigma)M=\operatorname{diag}(\sigma)\}\subset\mathbb{R}^{n\times n}.$ (34)
@@
If $s_{+}=0$ or $s_{-}=0$, this is the orthogonal group, otherwise it is the “indefinite” orthgonal group. The group $O(s_{+},s_{-})$ acts on $\tau_{r}^{-1}(S)\cap\mathcal{V}_{\sigma}$ according to $(M,(\sigma,W))\mapsto(\sigma,MW)$ since $W^{T}M^{T}\operatorname{diag}(\sigma)MW=W^{T}\operatorname{diag}(\sigma)W$. We claim that this action is transitive (any point in $\tau_{r}^{-1}(S)\cap\mathcal{V}_{\sigma}$ can be mapped to any other) and free (only the identity fixes any point). Indeed, if $W$ and $W^{\prime}$ are matrices in $\mathbb{R}^{r\times n}$ such that $W^{T}\operatorname{diag}(\sigma)W=W^{\prime T}\operatorname{diag}(\sigma)W^{\prime}=S$, then $Row(W)=Row(W^{\prime})=Row(S)$, and there exists $M\in GL(r,\mathbb{R})$ such that $W^{\prime}=MW$. Such $M$ belongs to $O(s_{+},s_{-})$ since $W^{T}M^{T}\operatorname{diag}(\sigma)MW=W^{T}\operatorname{diag}(\sigma)W$ implies $M^{T}\operatorname{diag}(\sigma)M=Id_{r}$ (by multiplying by the pseudoiverses of $W^{T}$ and $W$). This proves that the action is transitive. The action is also free since if $(\sigma,W)\in\tau_{r}^{-1}(S)$ then $rk(W)=r$ and $MW=W$ implies $M=Id$. The fact that the action is transitive and free means that the number of connected components of $\tau_{r}^{-1}(S)\cap\mathcal{V}_{\sigma}$ is the same as the number of connected components of $O(s_{+},s_{-})$. It is known that $O(s_{+},s_{-})$ has two connected components if $s_{+}=0$ or $s_{-}=0$, and four connected components otherwise [knapp2013lie]. In the former case, the two connected components correspond to the sign of the determinant, while in the latter case the components arise from the combinations of signs of the determinants of the two diagonal blocks of sizes $s_{+}\times s_{+}$ and $s_{-}\times s_{-}$. This proves cases $2$ and $3$.
@@
We now assume that $rk(S)<r$. We say that $S=W^{T}\operatorname{diag}(\alpha)W$ is a “minimal factorization” of $S$ if $\alpha$ has exactly ${\rm rk}(S)=s_{+}+s_{-}$ non-zero elements that are all $1$ or 1-1 and if $w_{i}=0$ for all $i$ such that $\alpha_{i}=0$. In this setting, $\alpha$ must have exactly $s_{+}$ elements that are 11 and $s_{-}$ elements that are 1-1 (Lemma 39). We will first prove that for every element $(\alpha,W)\in\tau_{r}^{-1}(S)$, we can find a path inside $\tau_{r}^{-1}(S)$ from $(\alpha,W)$ to a minimal factorization. Then we will show that any two minimal factorizations of $S$ are path connected in $\tau_{r}^{-1}(S)$.
@@
For the first claim, we observe as before that it is always possible to rescale $\alpha_{i}$ so that all of its elements are 1,11,-1 or 0. We thus need to show that we may transform $\alpha$ so it has exactly ${\rm rk}(S)$ non-zero elements. We prove this inductively arguing that if ${\rm rk}(S)<r$ and $W^{T}\operatorname{diag}(\alpha)W=S$, we can find a path inside $\tau_{r}^{-1}(S)$ from $(\alpha,W)$ to $(\alpha^{\prime},W^{\prime})$ so that $\alpha_{i}^{\prime}=w_{i}^{\prime}=0$ for some $i$. The same argument can then be repeated for the remaining $r^{\prime}=r-1$ vectors $w_{i}$ and elements $\alpha_{i}$ and iterated until $r^{\prime}={\rm rk}(S)$. If $\alpha_{i}=0$, then we can simply scale the corresponding $w_{i}$ and, similarly, if $w_{i}=0$ we can scale the corresponding $\alpha_{i}$. If $\alpha_{i}\neq 0$ and $w_{i}\neq 0$ for all $i$, then necessarily ${\rm rk}(W)<r$, so there exists $v\in\mathbb{R}^{r}$ such that $v^{T}W=0$. Assume that $\alpha$ has the first $\alpha_{+}$ elements equal to $1$, and the last $\alpha_{-}$ elements equal to 1-1 (with $\alpha_{+}+\alpha_{-}=r$). If $\alpha_{+}=0$ or $\alpha_{-}=0$, then we consider a path $(\alpha(t),W(t))$ where $\alpha(t)=\alpha$, $W(t)=M(t)W$ and $M(t)$ is a path in $SO(r)$ such that $M(0)=Id_{r}$ and the first row of $M(1)$ is equal to $v/\|v\|$ where $v^{T}W=0$ (this path exists because $SO(r)$ is connected). By construction the first row $M(1)W$ is zero, and we the can rescale the corresponding $\alpha_{1}$ to zero, as desired. If $\alpha_{+}>0$ and $\alpha_{-}>0$ then we apply essentially the same argument using the connected components of the indefinite orthogonal group $O(\alpha_{+},\alpha_{-})=\{M\in\mathbb{R}^{r\times r}\colon M^{T}\operatorname{diag}(\alpha)M=\operatorname{diag}(\alpha)\}$. However, we must first consider the special situation that our chosen $v\in\mathbb{R}^{r}\setminus\{0\}$ that satisfies $v^{T}W=0$ is such that $v^{T}\operatorname{diag}(\alpha)v=0$ (i.e., it is an isotropic vector). If this is the case, we construct a path as follows. Letting $y=\operatorname{diag}(\alpha)v$ and assuming that $y_{i}$ is a non-zero element of $y$ (note $y\neq 0$), we set $W(t)=W-(t/y_{i})yw_{i}^{T}$. Observing that $\operatorname{diag}(\alpha)y=v$ and that $y^{T}\operatorname{diag}(\alpha)y=v^{T}\operatorname{diag}(\alpha)^{3}v=0$ (because $\operatorname{diag}(\alpha)\operatorname{diag}(\alpha)=Id$), we have that
$(W-(t/y_{i})yw_{i}^{T})^{T}\operatorname{diag}(\alpha)(W-(t/y_{i})yw_{i}^{T})=W^{T}\operatorname{diag}(\alpha)W,$
@@
and the $i$-th row of $W-(1/y_{i})yw_{i}^{T}$ is 0, as desired. We finally assume that $v\in\mathbb{R}^{k}$ for which $v^{T}W=0$ is such that $v^{T}\operatorname{diag}(\alpha)v\neq 0$. We can normalize $v$ such that $v^{T}\operatorname{diag}(\alpha)v=1$ or $v^{T}\operatorname{diag}(\alpha)v=-1$. If $v^{T}\operatorname{diag}(\alpha)v=1$, then we can find a path $M(t)$ in $O(\alpha_{+},\alpha_{-})$ such that $M(0)=Id_{r}$ and the first row of $M(1)$ is equal to $v$ or v-v (this is true in light of the description of the connected components of $O(\alpha_{+},\alpha_{-})$). [^4] It is helpful to note that $M\in O(\alpha_{+},\alpha_{-})$ if and only if $M^{T}\in O(\alpha_{+},\alpha_{-})$. Indeed: $M^{T}\operatorname{diag}(\alpha)M=\operatorname{diag}(\alpha)\Leftrightarrow M^{T}\operatorname{diag}(\alpha)M\operatorname{diag}(\alpha)M^{T}=M^{T}\Leftrightarrow M\operatorname{diag}(\alpha)M^{T}=\operatorname{diag}(\alpha)$. Similarly, if $v^{T}\operatorname{diag}(\alpha)v=-1$, we can find a path $M(t)$ in $O(\alpha_{+},\alpha_{-})$ such that $M(0)=Id_{r}$ and the $(\alpha_{+})+1$-th row of $M(1)$ is equal to $v$ or v-v. Either way, we have that $(\alpha,W(t))=(\alpha,M(t)W)$ is a path in $\tau_{r}^{-1}(S)$ such that $W(0)=W$ and and $W(1)$ has a zero row. This concludes the proof of the existence of a path in $\tau_{r}^{-1}(S)$ from any $(\alpha,W)$ to a minimal factorization of $S$.